riesse922:

Juan used the expression 16 – 9 – 12 + 22 to find his profit for days 2 and 3. He rewrote the expression as 16 + (–9) + (–12) + 22. Juan can use the associative and commutative properties to rewrite the expression again. Explain why he had to use the additive inverse before he could use these properties.

riesse922:

hey i need help with this one and i was not get it

jimthompson5910:

We can add two numbers in any order, but we cannot subtract in any order Eg: 2+3 = 3+2 = 5 but 2-3 = -1 and 3-2 = 1 So the order doesn't matter with addition, but it does with subtraction When he wrote 16-9 as 16+(-9), this turned the subtraction problem into an addition one. After this point, he can swap the terms like so -9+16 and this is using the commutative property of addition

riesse922:

thank u

jimthompson5910:

no problem